OpenStudy (tanyatazf):

Let Z represent a standard normal random variable. P(Z>0) is equal to 0.0, 0.5, 0.45, or 0.9?

OpenStudy (princeharryyy):

what do u intend to ask?

OpenStudy (tanyatazf):

I don't know =[

OpenStudy (princeharryyy):

it's like this right. http://stattrek.com/probability-distributions/standard-normal.aspx

OpenStudy (nevillesnetwork):

which number is more than 0? The answer will be more than 0.

OpenStudy (princeharryyy):

No @NevillesNetwork It is from standard probability . It doesn't ask what will be z if z>0

OpenStudy (nevillesnetwork):

oh. XD oops

OpenStudy (princeharryyy):

they must have given u the table @tanyatazf just like in the link I gave you.

OpenStudy (princeharryyy):

all you need to do is find P(z>0) = 1 - P(z<0) You can find P(z<0) from the table they would have given you.

OpenStudy (tanyatazf):

They didn't. They just gave me the question as is. The answer was .50, I guessed. I don't know how it was found though.

OpenStudy (princeharryyy):

Than the table might be universal.

OpenStudy (princeharryyy):

but they have given an example in the link I gave you and they have done it in an example you can read that and find out how they have done it.

OpenStudy (tanyatazf):

Thank you so much for your help =]

OpenStudy (princeharryyy):

I found this calculator online. http://stattrek.com/online-calculator/normal.aspx

OpenStudy (princeharryyy):

All you have to do is put 0 on standard score. it will calculate you P(z< = 0) = 0.50 your answer P(z>0) = 1- P(z<0) = 1 -0.50 = 0.50 your answer.